# Loading and decay dynamics of an optical dipole trap using spatio-temporal Monte Carlo method

### v1

@@ II.1 Estimating the loss parameters
In this work, we have accounted for the non-uniform density of atoms in the dipole trap when applying the GP model. We have also considered the AC-stark shift of atomic energy levelsArora, Safronova, and Clark (2007) between $5S_{1/2}$ (F=2, $m_{F}=+2$) and $5P_{3/2}$ (F=3, $m_{F}=3$) while applying the GP-model in our simulations. A spatial Monte Carlo (MC) method is integrated in this model by randomly choosing atom pairs at various positions inside the trap as per Gaussian density distributions in axial and longitudinal directions. The effect of varying trap depth and distance between two atoms in an interacting pair is incorporated into $P_{re}$ and v(R,E) calculations. We eventually evaluate $K_{re}$. The evaluation of Eq. 2 has been done using Monte-Carlo integration with random number sampling of relative distance and energy.
@@
In the absence of near resonant MOT beams, the elastic evaporative collisions induce momentum transfer with energy of an atom sufficient to get kicked out of the dipole trap. They only lead to one atom loss from the trap. In our semiclassical model (Fig. 1), we incorporate quantum scattering due to low-energy s-wave collisions which are energy independent to approximate the effective cross-section of interaction ($\sigma_{el}$) van Kempen et al. (2002). The two body collision ($R_{el}$) can be approximated as,
@@
where $\sigma_{el}=8\pi a_{s}^{2}$ and $a_{s}$ ($\approx 5.3\,nm$) is the low energy s-wave scattering length van Kempen et al. (2002) for ${}^{87}Rb$ atoms, $v_{r}$ is average relative velocity of dipole trap atoms in the direction of approach and N is the number of atoms in dipole trap. For our system, the trap frequency in radial direction is $\approx 10$ kHz and in z-direction is $\approx 100$ Hz. Since, both these values are greater than estimated range of single-atom elastic collision rate. Thus the cold atom cloud system in consideration is in non-hydrodynamic regime and the collision rate scales as $N^{2}$. The effective volume $V_{eff}$ is given as,
@@
where $a_{r}\approx 6\mu m$ is the extent of the cloud size in transverse direction and $a_{z}\approx 436\mu m$ is the corresponding extent of the cloud in longitudinal direction of the dipole beam propagation. Finally, the two-body loss coefficient can be estimated accounting for the fraction of loss causing evaporative collisions. This is given as,
$\beta_{decay}=\frac{R_{ab}}{N^{2}}\frac{2}{\sqrt{\pi}}\eta e^{-\eta}$ (9)
where $\eta=\frac{U(r,z)}{k_{B}T_{OD}}$ dictates the probability of loss events. Here U(r,z) is the position dependent dipole trap potential. The Eq. 9 is evaluated by using spatial MC accounting for position dependent density and trap potential as discussed before for radiative escape. There is only a fraction of elastic collision events Ketterle and Druten (1996) within the effective range leading to loss of an atom.
Further, in absence of MOT beams, where we study decay dynamics in dipole trap, the contribution of dark ground-state hyperfine changing collisions Gensemer et al. (2000) to the overall loss rate can be neglected. This is because most atoms in our system are in the dark (F=1) ground state as the re-pumper beams are turned off slightly before the MOT beams. The hyperfine energy scale ($\approx 350\mu K$) is much higher than the available energy scale ($\approx 65\mu K$) of the trapped atoms in the ODT. Thus, this collision channel is effectively suppressed. The spin-changing collisions are also suppressed as there is no Zeeman splitting because magnetic fields are turned off during decay of dipole trap. Only the far-detuned dipole beam is present so no light assisted losses are possible and photo-association collisions are also suppressed heavily.
@@ III Experimental
[…]
@@
The dipole trap is formed by using the output of a master oscillator power amplifier (MOPA) system (TOPTICA TA PRO) operating at wavelength 811 nm. The laser beam was focused onto the UHV-MOT cloud using a 200 mm focal length lens mounted on a 3-axis translation stage. The trap beam has a $1/e^{2}$ radius of 12.9 $\mu m$ and the trap is loaded from the MOT in a duration of about 100 ms. The optical dipole trap (ODT) was loaded by switching-on the dipole laser beam on MOT cloud. The trapped cloud in ODT was imaged by absorption probe imaging technique. Using an appropriate image processing software, the absorption images were converted into optical density (OD) images. A fine straight line at the center of the image of atom cloud was observed at dipole laser power of $\approx$ 130 mW (corresponding to a trap depth of $\approx 77\mu K$). This showed the trapping of cold atoms in ODT as seen in Fig. 4. The lifetime study has been performed for 12.9 $\mu m$ trap after turning off the MOT beams. The single MOT beam intensity at the position of dipole trap center is around 17 $mW/cm^{2}$.
Figure 4: The absorption probe image of cold atom cloud in the dipole trap
@@ IV.1 Simulation results for loss parameters
The probability of radiative escape was calculated using the classical model developed in section II . The probability for radiative escape ($P_{re}$) after excitation at a certain pair distance (Fig. 5 A) within the atom cloud is found . The value is averaged over all pair kinetic energies sampled from Maxwell-Boltzmann distribution of pair kinetic energy (Eq. 5) and trap positions sampled from normal gaussian distributions taking density variation into account. The variation of $P_{re}$ is also shown with varying dipole trap depth (Fig. 5 B) and averaged over pair distance and pair energy distribution (Eq. 5) .
@@
We have also calculated the $K_{re}$ (shown in Fig. 6) using GP model (Eq. 3) integrated with spatial MC with varying trap depths (due to power as beam waist is kept constant). We see that as expected the collision rate decreases with increasing AC stark shift (or power) due to larger net detuning observed for excitation and reduced net absorption cross section. Similarly, we also see a decrease in the calculated collision rates with increase in detuning of MOT beams due to decrease in net absorption cross-section. We have also compared the calculated results (Fig. 7) for both the models of radiative escape discussed with varying power. The semi-classical model without pair-distribution function also shows a decrease in collision rate as power is increased. The loss coefficients can be extracted from these models by a factor of 2 multiplication.
@@ IV.2 Optical dipole trap loading and decay
The value of the estimated loading rate for the MOT parameters of density ($\approx\frac{10^{10}}{cm^{3}}$) and temperature ($\approx 50\mu K$) is around $R_{l}\approx 10^{6}$ atoms/s. The estimated value for $\gamma=0.25s^{-1}$ is known from the loss rate of MOT atoms. This serves as a decent approximation to initiate the temporal MC as the $\gamma$ is mainly dependent on background same for the traps. A more accurate estimate is not needed as the dynamics is dominated mainly by intra-trap losses. The effective volume of the cloud extracted from the OD images is $\approx 2.6\times 10^{-7}cm^{3}$. The estimated values of two body loss coefficients for fine-structure changing collisions ($\beta_{fs}\approx 2\times 10^{-4}s^{-1}$), bright ground state HF changing collisions ($\beta_{hf}\approx 1.9\times 10^{-4}s^{-1}$) and the radiative escape collisions ($\beta_{re}\approx 4.6\times 10^{-3}s^{-1}$) are calculated. This gives a collective two body loss coefficient $\beta_{LAC}\approx 5\times 10^{-3}s^{-1}$. However, if we use the estimate of two body loss coefficient ($\approx 8\times 10^{-3}s^{-1}$) obtained using average radiative escape probability model we get collective two body loss coefficient $\beta_{LAC}\approx 8.4\times 10^{-3}s^{-1}$.
The estimated parameters are used to initiate the MC data generation. The value of the final parameters obtained from our MC method (refer fig 7 A) applied on the experimental data are $R_{load}=6.7\times 10^{6}s^{-1}$, $\gamma=0.2s^{-1}$ and $\beta_{MC}=7.6\times 10^{-3}s^{-1}$. The value of the estimated two body loss coefficient is in good agreement with the experimentally observed value. The differential equation model Julienne and Vigué (1991) (Fig. 8B) with the inclusion of loading term has been used as well to fit the data as well and the loss coefficient obtained is $\beta_{DE}=(7.4\pm 0.7)\times 10^{-3}s^{-1}$. In all the earlier studies concerning cold $Rb^{87}$ atoms it has been observed that fine-structure changing collisions is a dominant loss channel as far as MOT is concerned. This observation of tremendous enhancement of radiative escape channel can be attributed to the very small trap depth of our dipole trap compared to MOT trap depths.
@@
For decay dynamics, the estimated value for $\gamma=0.25s^{-1}$ and that of $\beta^{{}^{\prime}}_{decay}=1.6\times 10^{-4}s^{-1}$ is obtained by the model of elastic evaporative loss discussed for a cold Van der Waals gas. The values given by the experimental data via Monte Carlo fitting (refer fig 7 B) and error minimization are $\gamma=(0.4\pm 0.15)s^{-1}$, $\beta=(1.8\pm 0.4)\times 10^{-4}s^{-1}$. The fitting via differential equation has been shown in same plot, suggesting a better fit via MC technique. The value of $\beta$ obtained from differential equation fitting is $(2.4\pm 0.2)\times 10^{-4}s^{-1}$. The semiclassical model seems to adequately capture most of the features of the collisions.

### v2

# Spatio-temporal Monte Carlo modeling of loading and decay dynamics in an optical dipole trap

@@ II.1 Estimating the loss parameters
In this work, we have accounted for the non-uniform density of atoms in the dipole trap when applying the GP model. We have also considered the AC-stark shift of atomic energy levelsArora, Safronova, and Clark (2007) between $5S_{1/2}$ (F=2, $m_{F}=+2$) and $5P_{3/2}$ (F=3, $m_{F}=3$) while applying the GP-model in our simulations. A spatial Monte Carlo (MC) method is integrated in this model by randomly choosing atom pairs at various positions inside the trap as per Gaussian density distributions in axial and longitudinal directions. The effect of varying trap depth and distance between two atoms in an interacting pair is incorporated into $P_{re}$ and v(R,E) calculations. We eventually evaluate $K_{re}$. The evaluation of Eq. 2 has been done over appropriate range of R and E values. The upper cutoff for R is decided based on comparison of dipole-dipole interaction energy with thermal energy. The upper cut-off for E is kept as twice the trap depth.
@@
In the absence of near resonant MOT beams, the intra-trap elastic evaporative collisions induce momentum transfer with energy of an atom sufficient to get kicked out of the dipole trap. They only lead to one atom loss from the trap. In our semiclassical model (Fig. 1), we incorporate quantum scattering due to low-energy s-wave collisions which are energy independent to approximate the effective cross-section of interaction ($\sigma_{el}$) van Kempen et al. (2002). The two body collision ($R_{el}$) can be approximated as,
@@
where $\sigma_{el}=8\pi a_{s}^{2}$ and $a_{s}$ ($\approx 5.3\,nm$) is the low energy s-wave scattering length van Kempen et al. (2002) for ${}^{87}Rb$ atoms, $v_{r}$ is average relative velocity of dipole trap atoms in the direction of approach and N is the number of atoms in dipole trap. For our system, the trap frequency in radial direction is $\approx 10$ kHz and in z-direction is $\approx 100$ Hz. Both these values are greater than estimated range of single-atom elastic collision rate ($\frac{N\sigma_{el}v_{r}}{V_{eff}}$). Thus the cold atom cloud system in consideration is in non-hydrodynamic regime and the collision rate scales as $N^{2}$. The effective volume $V_{eff}$ is given as,
@@
where $a_{r}\approx 6\mu m$ is the extent of the cloud size in transverse direction and $a_{z}\approx 436\mu m$ is the corresponding extent of the cloud in longitudinal direction of the dipole beam propagation. Similar to the LAC case we define an intra-trap elastic collision coefficient as $K_{el}=\frac{R_{el}}{N^{2}}$. Finally, the two-body loss coefficient can be estimated accounting for the fraction of loss causing evaporative collisions. This is given as,
$\beta_{dec}=K_{el}\frac{2}{\sqrt{\pi}}\eta e^{-\eta}$ (9)
where, $\eta=\frac{U_{dip}(r,z)}{k_{B}T_{OD}}$ dictates the probability of loss events. Here, $U_{dip}(r,z)$ is the position dependent dipole trap potential. The Eq. 9 is evaluated by using spatial MC accounting for position dependent density and trap potential as discussed before for radiative escape. There is only a fraction of elastic collision events Ketterle and Druten (1996) within the effective range leading to loss of an atom.
Further, in absence of MOT beams, where we study decay dynamics in dipole trap, the contribution of dark ground-state hyperfine changing collisions Gensemer et al. (2000) to the overall loss rate can be neglected. This is because most atoms in our system are in the dark (F=1) ground state as the re-pumper beams are turned off slightly before the MOT beams. The hyperfine energy scale ($\approx 350\mu K$) is much higher than the available energy scale ($\approx 65\mu K$) of the trapped atoms in the ODT. Thus, this collision channel is effectively suppressed. The spin-changing collisions are also suppressed as there is no Zeeman splitting because magnetic fields are turned off during decay of dipole trap. Only the far-detuned dipole beam is present so no light assisted losses are possible.
@@ III Experimental
We have experimentally studied the loading and decay dynamics of a dipole trap of beam waist 12.9 $\mu m$ ($\frac{1}{e^{2}}$ radius) . Our experiments (Fig. 3) have been performed using a single-chamber magneto-optical trap (MOT) for laser-cooled 87Rb atoms Bhardwaj et al. (2023) which has been used to load a single-beam optical dipole trap (ODT). The setup consists of a glass-cell ultra-high vacuum (UHV) chamber (UHV-MOT) at $\sim 1.5\times 10^{-10}$ Torr. Cooling beams for the MOT are generated using a high power tunable diode laser system (TOPTICA TA-PRO). The cooling laser is detuned by 15 MHz from the $35S_{1/2},F=2\rightarrow 5P_{3/2},F^{\prime}=$ 3 transition of 87Rb atom. The output of the laser system is expanded to form the UHV-MOT (six-beam configuration). The repumper laser beams are generated using an extended cavity diode laser system (TOPTICA DLC-PRO) which is stabilized to the transition the $F=1\rightarrow F^{\prime}=2$ of 87Rb atom. The quadrupole magnetic fields is provided by two coil pairs, and Rb vapor is supplied by resistively heated Rb-dispensers. With proper laser locking, magnetic field tuning, and beam alignment, we routinely load $\sim 5\times 10^{7}$ atoms into the UHV-MOT at a temperature of $\sim$50 $\mu$K after implementing molasses stage.
@@
The dipole trap is formed by using the output of a master oscillator power amplifier (MOPA) system (TOPTICA TA PRO) operating at wavelength 811 nm. The laser beam was focused onto the UHV-MOT cloud using a 200 mm focal length lens mounted on a 3-axis translation stage. The trap is loaded from the MOT in a duration of about 100 ms. The optical dipole trap (ODT) was loaded by switching-on the dipole laser beam on MOT cloud. The trapped cloud in ODT was imaged by absorption probe imaging technique. Using an appropriate image processing software, the absorption images were converted into optical density (OD) images. A fine straight line at the center of the image of atom cloud was observed at dipole laser power of $\approx$ 130 mW (corresponding to a trap depth of $\approx 77\mu K$). This showed the trapping of cold atoms in ODT as seen in Fig. 4. The lifetime study has been performed for 12.9 $\mu m$ trap after turning off the MOT beams. The single MOT beam intensity at the position of dipole trap center is around 17 $mW/cm^{2}$.
Figure 4: The experimental optical density image of the trapped atom cloud obtained by absorption imaging
@@ IV.1 Simulation results for loss parameters
The probability of radiative escape was calculated using the model developed in section II . In all these calculations, cloud volume $V_{eff}\approx 2.6\times 10^{-7}cm^{3}$ estimated based on temperature obtained from OD images and Eq. 8 was taken as constant. We also assumed a temperature roughly proportional to the trap depth Kuppens et al. (2000). The probability for radiative escape ($P_{re}$) after excitation at a certain pair distance (Fig. 5 A) within the atom cloud is found . The value is averaged over all pair kinetic energies sampled from Maxwell-Boltzmann distribution of pair kinetic energy (Eq. 5) and trap positions sampled from normal gaussian distributions taking density variation into account. The variation of $P_{re}$ is also shown with varying dipole trap depth (Fig. 5 B) and averaged over pair distance and pair energy distribution (Eq. 5) .
@@
We have also calculated the $K_{re}$ (shown in Fig. 6) using GP model (Eq. 3) integrated with spatial MC with varying trap depths (due to power as beam waist is kept constant). We see that as expected the collision rate decreases with increasing AC stark shift (or power) due to larger net detuning observed for excitation and reduced net absorption cross section. Similarly, we also see a decrease in the calculated collision rates with increase in detuning of MOT beams due to decrease in net absorption cross-section. We have also compared the calculated results (Fig. 7) for both the models of radiative escape discussed with varying power. The semi-classical model without pair-distribution function also shows a decrease in collision rate as power is increased. The inelastic loss coefficients are simply twice the collision coefficients as each collision effectively results in both atoms being lost.
In the absence of near-detuned MOT beams, we have simulated the elastic evaporative loss coefficient ($\beta_{dec}$) variation with the trap depth (Fig. 8). We assume a constant cloud volume ($V_{eff}$) for the trap depths considered and a temperature roughly proportional to trap depth as mentioned above for radiative escape. The loss increases as trap becomes tighter due to these elastic collisions. Each such collision only leads to a single atom being lost from the trap.
Figure 8: The coefficient of elastic evaporative collisions with the dipole trap depth for 12.9 micron beam waist is shown. The points are generated via spatial MC simulation.
@@ IV.2 Optical dipole trap loading and decay
The value of the estimated loading rate for the MOT parameters of density ($\approx\frac{10^{10}}{cm^{3}}$) and temperature ($\approx 50\mu K$) is around $R_{l}\approx 10^{6}$ atoms/s. The estimated value for $\gamma=0.25s^{-1}$ is known from the loss rate of MOT atoms. This serves as a decent approximation to initiate the temporal MC as the $\gamma$ is mainly dependent on background same for the traps. A more accurate estimate is not needed as the dynamics is dominated mainly by intra-trap losses. The estimated values of two body loss coefficients for fine-structure changing collisions ($\beta_{fs}\approx 2\times 10^{-4}s^{-1}$), bright ground state hyperfine changing collisions ($\beta_{hf}\approx 1.9\times 10^{-4}s^{-1}$) and the radiative escape collisions ($\beta_{re}\approx 4.6\times 10^{-3}s^{-1}$) are calculated. This gives a collective two body loss coefficient $\beta_{LAC}\approx 5\times 10^{-3}s^{-1}$. However, if we use the estimate of two body loss coefficient ($\approx 8\times 10^{-3}s^{-1}$) obtained using average radiative escape probability model we get collective two body loss coefficient $\beta_{LAC}\approx 8.4\times 10^{-3}s^{-1}$.
The estimated parameters are used to initiate the MC data generation. The value of the final parameters obtained from our MC method (Fig . 9A) applied on the experimental data are $R_{load}=6.7\times 10^{6}s^{-1}$, $\gamma=0.2s^{-1}$ and $\beta_{MC}=7.6\times 10^{-3}s^{-1}$. The value of the estimated two body loss coefficient is in good agreement with the experimentally observed value. The differential equation model Julienne and Vigué (1991) (Fig. 9A) with the inclusion of loading term has been used as well to fit the data as well and the loss coefficient obtained is $\beta_{DE}=(7.4\pm 0.7)\times 10^{-3}s^{-1}$. In all the earlier studies concerning cold $Rb^{87}$ atoms it has been observed that fine-structure changing collisions is a dominant loss channel as far as MOT is concerned. This observation of tremendous enhancement of radiative escape channel can be attributed to the very small trap depth of our dipole trap compared to MOT trap depths.
@@
For decay dynamics, the estimated value for $\gamma=0.25s^{-1}$ and that of $\beta^{{}^{\prime}}_{decay}=1.6\times 10^{-4}s^{-1}$ is obtained by the model of elastic evaporative loss discussed for a cold Van der Waals gas. The values given by the experimental data via Monte Carlo fitting (Fig. 9 B) and error minimization are $\gamma=(0.4\pm 0.15)s^{-1}$, $\beta=(1.8\pm 0.4)\times 10^{-4}s^{-1}$. The fitting via differential equation has been shown in same plot, suggesting a better fit via MC technique. The value of $\beta$ obtained from differential equation fitting is $(2.4\pm 0.2)\times 10^{-4}s^{-1}$. The semiclassical model seems to adequately capture most of the features of the collisions.